OpenStudy (anonymous):

My answer is A Which of the following is the solution to the equation 25(z + 4) = 125 ? (3 points) Answer z = 5.5 z = 3.5 z = -2.5 z = -0.5

OpenStudy (mathstudent55):

Are you sure you have the correct problem and solution choices?

OpenStudy (anonymous):

\[25^{z+4}=125\] \[5^{2(z+4)}=5^3\] \[2z+8=3\] solve for \(z\)

OpenStudy (anonymous):

yes, i'm sure

OpenStudy (anonymous):

so it is not A, you have to solve \[2z+8=3\]\[2z=-5\]\[z=-\frac{5}{2}=-2.5\]

OpenStudy (anonymous):

the gimmick is to rewrite so both sides have the same base

OpenStudy (anonymous):

ohhh.... i see what youre doing...i was way offf!!! ooops

OpenStudy (anonymous):

thank you. this helps loads

OpenStudy (mathstudent55):

I see from satelite that z + 4 is an exponent. Now the solution choices make sense.

OpenStudy (anonymous):

on the left you have \(25=5^2\) and on the right you have \(125=5^3\)

OpenStudy (mathstudent55):

You meant 25^(z + 4) = 125